# VideoWebArena: Evaluating Long Context Multimodal Agents with Video Understanding Web Tasks

### v1

@@ 5.2 Human Performance
To understand the level of human performance expected on the tasks within VideoWA, three authors attempted the tasks and provide intermediate answers for a random sample of each unique task template in the factual retention set. Further details on the human evaluation set can be found in Appendix A.3. They also tested 74 unique skill retention tasks, with each task having 2 separate humans attempt the task, one with a tutorial and one without. The humans performed actions and recorded steps using the VideoWA action space. Humans achieve a success rate of 73.9% on factual retention tasks while only taking an average of 6.4 steps per task (Table 5). Additionally, the intermediate intent and intent performance are linearly correlated while LLMs are not, citing a deficiency in the agentic abilities of these models. For skill retention tasks, human performance registers 93.1% on WebArena tasks and 88.6% in VisualWebArena with tutorials, and 82.6% and 72.7% without tutorials. We see an intuitive drop in human performance and efficiency without tutorials. In terms of both task success rates and average number of steps taken, video-capable LLM agents lag behind significantly, further emphasizing the need to improve agentic reasoning with video capacity in today’s models.
@@ 6 Discussion
We present VideoWebArena, a rigorous video-based agent benchmark that tests the agentic ability of long-context multimodal models. We define a task taxonomy of video-based agent tasks, utilizing a wide coverage of task types including skill retention and factual retention to create a comprehensive test bed for the setting of video agents. We provide 2021 tasks that are all video-based, along with 74 manually created videos. According to our experiments, the baseline agent does not perform well on most of tasks compared with human performance. There is still a long way in developing intelligent agents. For future work, it is important to analyze the failure cases explore better video agent architectures with different LLMs on this benchmark. We hope our environment and benchmark facilitate improvement and additional work on improving long-context multimodal agents.
Acknowledgements
We would like to thank Jing Yu Koh, Gabriel Sarch, Jennifer Hsia, Trevor Moreci, and Nicholas Chermak for providing valuable feedback on our project and paper.
@@ Acknowledgements
The authors are committed to making this work reproducible. Our code is open-sourced and available at https://github.com/ljang0/videowebarena. Our videos are also available through Google Drive and Youtube. Our data details are provided in Section 3 and our models and prompts are specified in Section 4.
@@ A.1 Video Creation Details
Three authors of the paper created 74 original video tutorials with audio narrating the actions taken. The three authors evenly divided the videos based off the site the video was based on. We based these 74 tutorials off of 1641 intents in WebArena and VisualWebArena. Each of these tasks are mapped to a video, creating a skill retention task. Each video creator manually checked all of the tasks that they were to create a tutorial for before, then made sure the functionality of the task was shown in the tutorial. We post these videos on Youtube at https://www.youtube.com/@webarenawarrior. We also provide them online at a Google Drive link: https://drive.google.com/file/d/1E1hM2jn1mj5q5d_j9mlx_n5FkbpZS1Yb/view?usp=drive_link.
@@ B.2 Long-Context Specific Failure Modes
Appendix C Results
C.1 Additional Results
@@
Appendix D Agent Prompts
D.1 Video Agent Task Prompt
@@
D.2 Video Frame Agent Task Prompt
@@
D.3 Video Summary Agent Task Prompt
@@
D.4 Video Agent Intermediate Task Prompt
@@
D.5 Video Frame Agent Intermediate Task Prompt
@@
D.6 Video Summary Agent Intermediate Task Prompt
@@
D.7 Video Summarization Prompt
@@
D.8 Video Frame Summarization Prompt

### v2

@@ 5.2 Human Performance
To understand the level of human performance expected on the tasks within VideoWA, three authors attempted the tasks and provide intermediate answers for a random sample of each unique task template in the factual retention set. Details on the human evaluation set can be found in Appendix A.3. They also tested 74 unique skill retention tasks, with each task having 2 separate humans attempt the task: one with a tutorial and one without. They performed actions and recorded steps using the VideoWA action space, achieving a success rate of 73.9% on factual retention tasks while only taking an average of 6.4 steps per task (Table 6). Additionally, the intermediate intent and intent performance are linearly correlated while LLMs are not, citing a deficiency in the agentic abilities of these models. For skill retention tasks, human performance registers 93.1% on WebArena tasks and 88.6% in VisualWebArena with tutorials, and 82.6% and 72.7% without tutorials. Naturally, we see a drop in human performance and efficiency without tutorials. In terms of both task success rates and average number of steps taken, video-capable LLM agents lag behind significantly, further emphasizing the need to improve agentic reasoning with video capacity in today’s models.
@@ 6 Discussion
We present VideoWebArena, a video-based agent benchmark that tests the video capability of long-context multimodal models through an agentic lens. We define a task taxonomy of video-based agent tasks, utilizing a wide coverage of task types including skill retention and factual retention to create a comprehensive test bed for the setting of video agents. We provide 2,021 tasks that are all video-based, along with 74 manually created videos.
From our experiments, we see that introducing video can result in “noise” on skill-retention tasks, causing the agent to fail at following action generating instructions. For instance, the agent would return try to click using the name of the element rather than the ID number. Another issue was where the agent would generate multi-action responses when the prompt explicitly says to generate one action. From analyzing agent trajectories, we see that agents show inconsistencies in both their reasoning and planning, such as getting stuck in a loop and repeating the same action/step even when the observed state or information does not change. We leave details to Appendix B. We hypothesize that further improvements can result from better distillation of multimodal chain of thought reasoning in the presence of an increasing number of modalities or having the agent learn which modalities are the most relevant for certain tasks over others.
Overall, the baseline agents do not perform well on most tasks especially relative to human performance—there is still a long ways ahead in developing intelligent agents capable of effectively utilizing and understanding video in completing regular tasks. For future work, we believe further analysis of failure modes from different video agent architectures on our benchmark may be key. We hope our environment/benchmark can help facilitate improvement and additional work on improving long-context multimodal agents.
@@
The authors are committed to making this work reproducible. Our code is will be open-sourced and available on Github. An anonymized version of our code base can be found at https://anonymous.4open.science/r/videowebarena-236E/README.md. Our videos will also be made available through Google Drive and YouTube. Our data details are provided in Section 3 and our models and prompts are specified in Section 4.
@@
We would also like to be cognizant of and make aware the possible biases that are introduced in the video tutorials that we have created. Our main focus is studying the ability of agents to make use of video information to accomplish tasks in dynamic settings. As the issues of multi-modal video-capable agent performance and in-context video understanding pose huge obstacles, we made this the main focus of our video and benchmark creation. As all the videos are in English and created by graduate students in an American university, our videos may unintentionally contain possible linguistic and cultural biases. To help mitigate this, we plan on adding Korean/Chinese/Portugese versions of the video dataset as well. Given the significant time and labor required to implement new websites and the main focus of our paper, adding non-English websites unfortunately is beyond the scope of this paper.
We have, however, made it so that our benchmark and its videos can be easily accessed and spun-up by the open-source community to test/study potential biases. We have intentionally left room for updates and extensions to the dataset to incorporate additional underrepresented content or address gaps identified by the community. Our work and its documentation make it very easy for users and the open-source community to create their own tasks customized with their own videos, websites, etc. that reflect specific variations of interest that may be underrepresented here. Existing websites can be used to create new tasks. We plan on monitoring our benchmark’s release and usage in the open-source community to address or mitigate possible biases as they arise or are made aware to us through changes in our benchmark or otherwise.
@@ A.1 Video Creation Details
Three authors of the paper created 74 original video tutorials with audio narrating the actions taken. The three authors evenly divided the videos based off the site the video was based on. We based these 74 tutorials off of 1641 intents in WebArena and VisualWebArena. Each of these tasks are mapped to a video, creating a skill retention task. Each video creator manually checked all of the tasks that they were to create a tutorial for before, then made sure the functionality of the task was shown in the tutorial. We post these videos on YouTube at https://www.youtube.com/@webarenawarrior. We also provide them online at a Google Drive link: https://drive.google.com/file/d/17DwmsM7KzBWyz1BN1aq7NHDvgcTIrCgx/view?usp=drive_link.
@@ B.2 Long-Context Specific Failure Modes
Appendix C Task Creation Documentation
C.1 Example Task JSON
We provide an example task below in their JSON file format.
"sites": [
"classifieds"
"task_id": 4,
"require_login": true,
"storage_state": "./.auth/classifieds_state.json",
"start_url": "__CLASSIFIEDS__",
"geolocation": null,
"intent_template": "Can you take me to {description}?",
"intent": "Can you take me to the cheapest primary red boat for sale in the video?",
"intermediate_intent": "What was the name of the post for the cheapest primary red boat for sale in the video?",
"video": "search_sort_comment",
"instantiation_dict": {
"description": "the cheapest primary red boat for sale in the video"
"require_reset": false,
"intermediate_eval": {
"eval_types": [
"string_match"
"reference_answers": {
"must_include": [
"1986 nova xl 23ft"
"reference_url": "",
"program_html": [],
"string_note": "",
"reference_answer_raw_annotation": ""
"eval": {
"eval_types": [
"url_match"
"reference_answers": null,
"reference_url": "__CLASSIFIEDS__/index.php?page=item&id=49894",
"program_html": [],
"url_note": "EXACT"
"intermediate_difficulty": "hard",
"overall_difficulty": "medium",
"visual_perception": true,
"audio_perception": false,
"full_video_understanding": true,
"temporal_reasoning": false,
"comments": ""
C.2 Task Creation Documentation
We used the following instructions to task creators. These can be used to create new tasks. Additionally, you can use any video as input to your LLM agent - our evaluation suite simply chooses from our own 74 manually made tutorials.
In order to create a task, choose or make a video of your own to reference to in order to create a task. Each task should be video-grounded, such that the task can only be completed after watching the video. For each new task, you need to fill the following values in a json task file:
  1. 1.
    Sites: The site(s) used in the task.
  2. 2.
    Start URL: The url the task starts on.
  3. 3.
    Video: The video name associated with the task.
  4. 4.
    Task ID: A unique ID for the task.
  5. 5.
    Intent: The prompt for the video-based task.
  6. 6.
    Intent Template: The task template specific to the intent with default parameters.
  7. 7.
    Intermediate Eval: Evaluates if your Vision-Language Model (VLM) can extract the necessary information from the video through a task-based VQA. More documentation on existing evaluators can be found in the evaluators.py file in our repository.
  8. 8.
    Eval: The evaluation suite to test whether the agentic task was completed successfully. More documentation on existing evaluators can be found in the evaluators.py file in our repository.
  9. 9.
    Intermediate Difficulty: The difficulty level of the intermediate VQA intent, distributed between easy, medium, and hard. The video difficulty ratings signify the complexity of returning the correct answer for a given task’s intermediate intent. Easy tasks require returning one piece of information and can be solved with less than 3 frames, medium tasks require returning 2 to 3 things and can be solved with less than half the video, and hard tasks require returning more than 3 things and require watching more than half the video.
  10. 10.
    Overall Difficulty: The overall difficulty level of the task, distributed between easy, medium, and hard. The agentic difficulty for each task signifies the complexity of the action sequence needed to complete an intent successfully. For agentic difficulty, we classify a task as easy if it can be completed in 1-3 steps, medium if it can be completed in 4-9 steps, and hard if it can be completed in more than 9 steps
  11. 11.
    Visual Perception: Whether visual perception of the video input is required. True or False.
  12. 12.
    Audio Perception: Whether audio reasoning of the video input is required. True or False.
  13. 13.
    Full Video Understanding: Whether full video comprehension of the video input is necessary. True or False.
  14. 14.
    Temporal Reasoning: Whether reasoning across different time points in the video input is required. True or False.
Appendix D Results
D.1 Additional Results
@@
Appendix E Agent Prompts
E.1 Video Agent Task Prompt
@@
E.2 Video Frame Agent Task Prompt
@@
E.3 Video Summary Agent Task Prompt
@@
E.4 Video Agent Intermediate Task Prompt
@@
E.5 Video Frame Agent Intermediate Task Prompt
@@
E.6 Video Summary Agent Intermediate Task Prompt
@@
E.7 Video Summarization Prompt
@@
E.8 Video Frame Summarization Prompt